OpenStudy (selena2345):

At A Typical National Football League Game the ratio of males to females in attendance is 3:2.There are 12,000 female spectators.What is the ratio of the number of male spectators to the total number of spectators?

OpenStudy (selena2345):

Hello Diyadiya where are you

OpenStudy (diyadiya):

Okay ,so number of female spectators =12000 \[\frac{Male}{Female}= \frac{3}{2}\] Let the number of Male spectators be x \[\frac{x}{12000}= \frac{3}{2}\]Multiply both sides by 12000\[\frac{x}{12000} \times 12000 = \frac{3}{2} \times 12000\]\[\frac{x}{\cancel{12000}} \times \cancel{12000} = \frac{3}{2} \times 12000\]\[x= \frac{3}{2} \times 12000\]\[x= 18000\]So we found out number of male spectators Now we have to find the ratio of number of male spectators to the total number of spectators Male spectators =18000 Total number of Spectators =18000+12000 = ? Now,\[\frac{male \ spectators}{Total\ spectators }= ?\] Can you find ?